OpenStudy (anonymous):

michele If f(x) = log2 (x + 4), what is f^–1(3)?

OpenStudy (anonymous):

@Michele_Laino 8 is answer?

OpenStudy (jessie13):

Are you trying to find out what x =?

OpenStudy (anonymous):

what is f^–1(3)

OpenStudy (michele_laino):

here we have: \[\Large {f^{ - 1}}\left( x \right) = {2^x} - 4\] so: \[\Large {f^{ - 1}}\left( 3 \right) = {2^3} - 4 = ...?\] more explanation: we can write: \[\Large f\left( {{f^{ - 1}}\left( x \right)} \right) = {\log _2}\left( {{2^x} - 4 + 4} \right) = {\log _2}\left( {{2^x}} \right) = x\] here's why the inverse function is: \[\Large {f^{ - 1}}\left( x \right) = {2^x} - 4\]

OpenStudy (anonymous):

right michele? 8 is right?

OpenStudy (michele_laino):

\[{f^{ - 1}}\left( 3 \right) = {2^3} - 4 = 8 - 4 = ...?\]

OpenStudy (anonymous):

4

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thank michele!

OpenStudy (michele_laino):

thanks! @sallyfield888